OpenStudy (samhang):

Integrate the following by Integration by partial fraction (y^2+4)/[y(3y^2+4y-4)

OpenStudy (jiteshmeghwal9):

\[\frac{A}{y}+\frac{By+c}{3y^2+4y-4}=\frac{(y^2+4)}{y(3y^2+4y-4)}\]\[\implies A(3y^2+4y-4)+(By+c)(y)=y^2+4\]\[\implies (3A+B) y^2+(4A+c)y-4A=y^2+4\]comparing we get 3A+B=1 4A+c=0 -4A=4 =>A=-1,c=4,B=4

OpenStudy (jiteshmeghwal9):

so ur equation reduces to \[\int\limits \left( \frac{-1}{y}+\frac{4y+4}{3y^2+4y-4} \right)dy\]

OpenStudy (jiteshmeghwal9):

solve it further & tell the result

OpenStudy (mathmale):

Hint: The first (of three) integrals is\[-\int\limits_{}^{}\frac{ dy }{ y }\] Can you evaluate this?

OpenStudy (samhang):

yes... \[-\ln \left| y \right|+C\]

OpenStudy (samhang):

result \[1/2\ln \left| y+2 \right|+5/6\ln \left| 3y-2 \right|-\ln \left| y \right|+C\]

OpenStudy (samhang):

I am no sure about but thanks lot guys for great help....